OpenStudy (anonymous):

find the mgf of f(y)=.5e^(-abs(y)).

OpenStudy (anonymous):

I know m(t) = E(e^ty)

OpenStudy (anonymous):

which is = to the integral of (e^ty)f(y)dy

OpenStudy (anonymous):

but I'm having trouble solving

OpenStudy (anonymous):

What's the support of the pdf?

OpenStudy (anonymous):

I suppose we're to assume it's all real \(y\)? \[m(t)=E\left[e^{ty}\right]=\frac{1}{2}\int_{-\infty}^\infty e^{ty}e^{-|y|}~dy\] Let's consider that absolute-valued exponent. When \(y<0\), you have \(|y|=-y\), and when \(y\ge0\), you have \(|y|=y\). So, split up the integral at 0 and you get \[\frac{1}{2}\left(\int_{-\infty}^0e^{ty}e^{-(-y)}~dy+\int_0^\infty e^{ty}e^{-y}~dy\right)\]

OpenStudy (anonymous):

pdf support is -infinity to +infinity

OpenStudy (anonymous):

Ah, good, on the right track then. \[\frac{1}{2}\left(\int_{-\infty}^0e^{ty}e^{y}~dy+\int_0^\infty e^{ty}e^{-y}~dy\right)\] \[\frac{1}{2}\left(\int_{-\infty}^0e^{y(t+1)}~dy+\int_0^\infty e^{y(t-1)}~dy\right)\] \[\frac{1}{2}\left(\frac{1}{t+1}\left[e^{y(t+1)}\right]_{-\infty}^0+\frac{1}{t-1}\left[e^{y(t-1)}\right]_0^\infty \right)\]

OpenStudy (anonymous):

why is the first part t+1 and the second t-1?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles Because \(e^{a}e^{b}=e^{a+b}\). In the first integral, the exponents are \(ty\) and \(y\), so \(e^{ty}e^{y}=e^{ty+y}=\color{red}{e^{y(t+1)}}\). The reasoning applies to other integrand's exponent. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

right. I understand the exponents and adding them and their rules. I'm more confused by the difference between the + and -.

OpenStudy (anonymous):

The plus and minus come as a result of the absolute value definition, which I used in my second post.

OpenStudy (anonymous):

but the |y| can never =-y which is what you wrote? or am i missunderstanding your post?

OpenStudy (anonymous):

Regardless of \(y\)'s sign, \(|y|\) is positive, yes, but \(|y|=y\) is not true if \(y\) is negative. For example, \(|-2|\not=-2\), since \(|-2|=2\). The minus sign in front of the \(y\) "corrects" for this. In this case, \(y=-2\), and so \(-y=2=|-2|=|y|\).

OpenStudy (anonymous):

Does that integral converge?

OpenStudy (anonymous):

That's what I was afraid of @wio. Checking with Wolfram, and looks like we need to set some restrictions on \(t\).

OpenStudy (anonymous):

okay. I think i understand

OpenStudy (anonymous):

but i'm still not sure how to continue

OpenStudy (anonymous):

Hey, neither do I! I still haven't covered continuous random variables, so I had to refer to my textbook :)

OpenStudy (anonymous):

I hate stats! :(

OpenStudy (anonymous):

The moment generating function really don't work here.

OpenStudy (anonymous):

but that's the question i was asked to do. :/

OpenStudy (anonymous):

I have to get some sleep, if you guys can help me out any further please let me know. Thanks for the help!